# Investigating Coverage Criteria in Large Language Models: An In-Depth Study Through Jailbreak Attacks

### v1

@@ III-C Standards for Coverage Evaluation
To evaluate the performance of different coverage criteria on LLMs, we propose three fundamental requirements inspired by traditional coverage criteria standards [10, 11, 12, 18] that focus on diversity assessment, fault detection, and generalization capabilities:
Requirement 1: Accurate Redundant Test Identification. Traditional coverage criteria must accurately reflect diversity within the test suite [12], meaning they should be insensitive to redundant tests. For LLMs, paraphrased sentences that share the same meaning as the original sentences and trigger the same or similar responses (indicating they activate the same “logic” of the LLM) should be regarded as redundant test cases. Therefore, an effective coverage criterion should not be overly sensitive to these synonymous queries.
Requirement 2: Sensitivity to Attack Queries. Traditional coverage demonstrates high discernibility for the abnormal behaviors caused by fault inputs, thereby assessing whether the test suite sufficiently explores unknown states of the model [18]. For LLMs, attack queries often lead to abnormal behaviors (jailbreak) to generate harmful content. Thus, an effective coverage criterion should be particularly sensitive to attack queries that can trigger abnormal behaviors.
Requirement 3: Stability and Generalization Ability. An effective coverage criterion should be robust to variations in LLM architecture and size, showing consistent effectiveness across different models. Given the rapid evolution of LLMs, a coverage criterion must have strong generalization capabilities to accurately guide model testing, regardless of the specific model evaluated. This requirement combines the generalization principles of traditional coverage criteria, emphasizing consistency and applicability across different models.
@@ IV-A1 Models
In this study, we comprehensively evaluate four well-known open-source LLMs, which vary significantly in size, architecture, and origin. These models include OPT-125M [23], Llama-2-7B-Chat [19], Pythia-12B [24], and Gemma-2-27B-it [25]. This selection covers a broad spectrum of model characteristics, ensuring an adequate observation across different dimensions.
@@ IV-A2 Test Suite Construction
To evaluate against Requirement 3, we analyze the coverage criterion’s performance across different models using the same test suite. Additionally, we construct test suites $S_{NM}$ and $S_{RM}$ that include rejected queries (which models tend to reject) to observe their impact on coverage. We expect a moderate coverage increase from adding these rejected queries, as they are distinct from normal queries but consistently rejected by the model. However, this outcome depends on the model’s security alignment and is not considered a requirement for evaluating the coverage criterion.
Therefore, to construct our test suites as listed in Table II, we select the following two widely-used datasets and create a complementary dataset on our own:
@@
Synonymous Query Dataset: To construct synonymous queries, we used GPT-4 to generate corresponding synonymous paraphrases for the first 500 queries from the Alpaca-gpt4 dataset, e.g., “What is the capital of France?” and “Name the capital city of France.”. This complementary dataset serves as the source of synonymous queries for the test suites.
TABLE II: Distribution of test suites across datasets
@@ IV-A3 Coverage Criteria Settings
NC requires an activation threshold parameter $T$ to determine whether a neuron is activated. Due to the different size and activation functions among OPT-125M, Llama-2-7B-Chat, Pythia-12B, and Gemma-2-27B-it, which significantly affect the distribution of neuron activations, we empirically set $T$ to 0.1, 0.25, 0.75, and 50 for the best performance, respectively.
TKNC requires a parameter $K$ to determine the number of top neurons selected. For all models, we set $K$ to 10.
TKNP is similar to TKNC. However, our experiments show that due to the complexity of LLMs, setting $K$ too high results in each new input forming a new pattern. Therefore, we set $K$ to 1.
@@ IV-B RQ1: Evaluating Coverage Criteria Effectiveness in LLMs
TABLE III: Coverage results for each criterion on different test suites. On $S_{N}$, the original coverage results are displayed. While for other test suites, the change rates relative to $S_{N}$ are demonstrated.
@@
To address RQ1, we evaluate the performance of the coverage criteria across different test suites and models to determine if it meets evaluation requirements. Table III presents the neuron coverage measured at the last query token across the Attention and MLP layers of each block. In RQ2 and RQ3, we will further investigate the impacts of different network layers and tokens on the coverage evaluation results.
@@
Evaluation against Requirement 2. Requirement 2 mandates that the coverage criteria be sensitive to attack queries. The average coverage growth on the Attention layer for NC, TKNC, TKNP, TFC, and NLC records at 15.23%, 17.37%, 32.99%, 45.88%, and 44.03% respectively, comparing $S_{NJ}$ to $S_{N}$. All five criteria achieve significant growth, with TFC and NLC showing the largest increases. The performance on the MLP layer, however, presents a slight variation, with coverage growth for the five criteria at 2.96%, 15.17%, 48.83%, 35.65%, and 3.92%, respectively. Coverage growth for NC and NLC is notably lower compared to their performance on the Attention layer, while TKNC, TKNP, and TFC continue to perform well. This indicates that all five criteria are capable of recognizing attack queries. However, performance varies across different model layers. We will discuss these layer-level differences in detail in RQ2. Due to Pythia-12B’s low safety alignment, questions that should be refused might still output malicious content, resulting in a higher coverage rate on $S_{NM}$, especially on the MLP layer where TFC and NLC coverage rates exceed those on $S_{NJ}$.
@@
Answer to RQ1: NC and TKNC are the most effective for LLMs. They consistently perform well in accurately identifying synonymous queries and detecting attack queries, while also maintaining stability and generalization across different model architectures and sizes.
@@ IV-C RQ2: Analyzing Model Layer-wise Contributions to Coverage in LLMs
Attention Layer versus MLP Layer: We first evaluate the performance of Attention and MLP layers from an overall model perspective. We use NC and TKNC, identified as effective in RQ1, as our research criteria and calculate RCG based on the data in Table III. For NC, the RCG values for Attention layers in the four models are 6.00%, 15.67%, 19.28%, and 7.35%, respectively. Correspondingly, the RCG values for MLP layers are 4.29%, 2.16%, 0.76%, and 0.89%. It is evident that under the same criteria and models, NC is more effective in Attention layers than in MLP layers.
Similarly, for TKNC, the RCG values for Attention layers in the four models are 6.10%, 11.65%, 10.46%, and 11.32%, while the MLP layers show values of 3.43%, 21.09%, 6.99%, and 3.67%. Except for Llama-2-7B-Chat, the effectiveness of Attention layers remains superior to that of MLP layers. Additionally, as shown in Table III, the $S_{NJ}$ coverage growth rates are highest for all criteria in Attention layers. In contrast, in MLP layers, higher $S_{NM}$ growth rates are observed in NC for the Llama-2-7B-Chat model, and in TFC and NLC for the Pythia-12B model.
@@
Figure 4: The RCG results based on NC and TKNC for different blocks of the four target LLMs. OPT-125M contains 12 blocks, Llama-2-7B-Chat contains 32 blocks, Pythia-12B contains 36 blocks, and Gemma-2-27B-it contains 46 blocks.
@@ IV-D RQ3: Investigating Token-level Impacts on Coverage Analysis in LLMs
Figure 5: The RCG results calculated based on NC and TKNC for different tokens in the target LLMs. Each model generates 10 tokens for each query.
To investigate RQ3, we expanded on the previous experiments by generating 10 additional tokens for each test suite, starting from the last token of the query  [^1] For experimental stability, we select the next token with the highest probability.. We explore how different tokens impact coverage analysis, assuming they may lead to varying coverage behaviors and reveal more comprehensive insights into model behavior. Our key question is: “How does generating additional tokens affect diversity assessment among test suites in LLM testing?”
Our further goal is to identify the optimal token positions for LLM testing to reduce computational costs and achieve efficient testing results. These points are crucial for determining the most effective moments to measure coverage, thereby optimizing the balance between testing and resource expenditure. After generating each new token, we calculated the corresponding coverage rates for the model. This process allowed us to examine how the coverage evolves as the model generates more tokens beyond the initial query. Similarly, we analyzed the impact of these tokens on coverage analysis through the RCG quantified by NC and TKNC. The experimental results are shown in Figure 5.
@@
V Application for Real-time Jailbreak Detection
Our empirical study shows that NC is an effective coverage criterion, and the attention layers and the last query token are more sensitive to different test suites. This indicates that the number of activated neurons (the feature used by NC) can effectively differentiate between normal queries and jailbreak attacks on the LLMs, inspiring us to design a jailbreak attack detection method utilizing this feature. Particularly, we achieve the detection by training a Multilayer Perceptron (MLP) model, using the number of activated neurons as input, to output a binary result indicating whether a query to LLMs could cause a jailbreak response. The dimension of input to the classifier model (i.e., the number of activated neurons) varies with the LLMs’ architecture. For example, OPT-125M has 12 blocks. We construct a feature vector $(l_{0},l_{1},\ldots,l_{11})$, consisting of exactly 12 features representing the number of activated neurons in the attention layers of these 12 blocks, as the input to the corresponding classifier. For each classifier, they uniformly include four hidden layers with 256, 2048, 512, and 128 neurons, respectively.
To validate the effectiveness of this coverage-criteria-based classifier, we design the following experiments on OPT-125M, Llama-2-7B-Chat, and Pythia-12B. For more details, please refer to our website [22].
V-A Effectiveness Comparison
Dataset: To train and test the classifier, we use four datasets. From Alpaca-gpt4, we randomly select 2500 queries for training and 500 for testing, representing normal queries to LLMs. Similarly, from JailBreakV-28k, we select 2500 attack queries for training and 500 for testing, representing attack queries that lead the model to output malicious content. The TruthfulQA dataset, containing 817 common life questions prone to errors, is used to verify how well the classifier generalizes to normal queries. For adversarial attack queries, we use GCG to generate adversarial attack suffixes for 200 rejected queries in AdvBench based on Llama-2-7B-Chat  [^2] The GCG repository is implemented only on Vicuna-7B and Llama-2-7B-Chat, but their high transferability allows application to other models.. This dataset is used to validate how well the classifier generalizes to attack queries.
@@ V-A Effectiveness Comparison
Baseline: To demonstrate the effectiveness of the classifier, we select the most widely used perplexity filter [27] and the state-of-the-art method, PARDEN [28], as baselines for comparison. We follow the settings in the  [28], where the threshold for the perplexity filter is set to the maximum perplexity of malicious queries in AdvBench, with a window size of 10. The threshold for PARDEN is set to 0.6, with a window size of 30.
TABLE IV: Experimental Results of the Classifier Constructed Based on NC

### v3

@@ III-C Standards for Coverage Evaluation
To evaluate the performance of different coverage criteria on LLMs, we propose three fundamental requirements inspired by traditional coverage criteria standards [14, 15, 16, 21, 20], focusing on functional diversity, defect-detection capability, and generalization. Due to the wide and unconstrained output space of LLMs, we define these requirements from the perspective of the jailbreak safety problem, extending traditional standards to address the unique challenges in LLM testing.
Requirement 1: Accurate Redundant Test Identification. Aligned with the emphasis on functional diversity in traditional DNN testing, an effective coverage criterion for LLMs should accurately reflect diversity within the test suite and be insensitive to redundant tests. For example, in traditional image classification, two similar images (non-adversarial) are often considered redundant test cases and do not contribute to the diversity of the test suite [16, 20]. In the context of LLMs, paraphrased queries that convey the same meaning and trigger similar responses—thus activating the same “logic”—should also be be regarded as redundant. This ensures that the coverage metric genuinely represents the breadth of the model’s explored behaviors without being inflated by synonymous inputs.
Requirement 2: Sensitivity to Attack Queries. Building upon the defect-detection capability emphasized in traditional coverage criteria, we extend this concept under the jailbreak perspective for LLMs. For example, in traditional DNN models, adversarial attacks often lead to incorrect classifications, which are considered faults in the model [16]. Similarly, in the context of LLMs, we regard abnormal behaviors caused by jailbreak attacks as faults. An effective coverage criterion should be particularly sensitive to attack queries that induce abnormal behaviors, such as generating harmful or disallowed content. By focusing on the model’s responses to malicious inputs, the coverage metric assesses the model’s robustness and its ability to handle security threats. This extension addresses the unique safety concerns associated with LLMs’ expansive output space while maintaining correspondence with traditional defect-detection objectives.
Requirement 3: Stability and Generalization Ability. Consistent with the generalization principles valued in traditional coverage standards, we adapt this requirement to the diverse landscape of LLMs. An effective coverage criterion should exhibit stability and strong generalization capabilities, providing consistent guidance for model testing regardless of specific model variations. Given the rapid evolution and diversity of LLMs, this ensures that the coverage metric remains effective across different models, paralleling traditional standards while addressing the specific challenges in LLM testing.
@@ IV-A1 Models
In this study, we comprehensively evaluate four well-known open-source LLMs, which vary significantly in size, architecture, and origin. These models include OPT-125M [47], Llama-2-7B-Chat [43], Pythia-12B [48], and Gemma-2-27B-it [49]. Note that we include OPT-125M and Pythia-12B, which are non-safety-aligned, to ensure comprehensive observations.
@@ IV-A2 Test Suite Construction
To evaluate against Requirement 3, we analyze the coverage criterion’s performance across different models using the same test suite.
Additionally, we construct test suites $S_{NM}$ and $S_{RM}$ that include rejected queries (which models tend to reject) to observe their impact on coverage. We expect a moderate coverage increase from adding these rejected queries, as they are distinct from normal queries but consistently rejected by the model. However, this outcome depends on the model’s security alignment and is not considered a requirement for evaluating the coverage criterion.
Therefore, to construct our test suites as listed in Table III, we select the following two widely-used datasets and create a complementary dataset on our own:
@@
Synonymous Query Dataset: To construct synonymous queries, we use GPT-4 to generate corresponding synonymous paraphrases for the first 500 queries from the Alpaca-gpt4 dataset and manually verify that the synonymous queries produce outputs similar to the original queries (e.g., “What is the capital of France?” and “Name the capital city of France.” both with the answer “Paris”). This complementary dataset serves as the source of synonymous queries for the test suites.
TABLE III: Distribution of Test Suites Across Datasets
@@ IV-A3 Coverage Criteria Settings
NC requires an activation threshold parameter $T$ to determine whether a neuron is activated. Due to the different size and activation functions among OPT-125M, Llama-2-7B-Chat, Pythia-12B, and Gemma-2-27B-it, which significantly affect the distribution of neuron activations, we empirically set $T$ to 0.1, 0.25, 0.75, and 50, respectively.
TKNC requires a parameter $T$ to determine the number of top neurons selected. For all models, we set $T$ to 10.
TKNP is similar to TKNC. However, our experiments show that due to the complexity of LLMs, setting $T$ too high results in each new input forming a new pattern. Therefore, we set $T$ to 1.
@@ IV-B RQ1: Evaluating Coverage Criteria Effectiveness in LLMs
TABLE IV: Coverage Results for Each Criterion on Different Test Suites: On $S_{N}$, the Original Coverage Results Are Displayed, While for Other Test Suites, the Change Rates Relative to $S_{N}$ Are Demonstrated
@@
To address RQ1, we evaluate the performance of the coverage criteria across different test suites and models to determine if it meets evaluation requirements. Table IV presents the neuron coverage measured at the last query token across the Attention and MLP layers of each block. In RQ2 and RQ3, we will further investigate the impacts of different network layers and tokens on the coverage evaluation results.
@@
Evaluation against Requirement 2. Requirement 2 mandates that the coverage criteria be sensitive to attack queries. The average coverage growth on the Attention layer for NC, TKNC, TKNP, TFC, and NLC records at 15.23%, 17.37%, 32.99%, 45.88%, and 44.03% respectively, comparing $S_{NJ}$ to $S_{N}$. All five criteria achieve significant growth, with TFC and NLC showing the largest increases. The performance on the MLP layer, however, presents a slight variation, with coverage growth for the five criteria at 2.96%, 15.17%, 48.83%, 35.65%, and 3.92%, respectively. Coverage growth for NC and NLC is notably lower compared to their performance on the Attention layer, while TKNC, TKNP, and TFC continue to perform well. This indicates that all five criteria are capable of recognizing attack queries. However, performance varies across different model layers. We will discuss these layer-level differences in detail in RQ2.
Furthermore, we observe significant coverage differences between normal ($S_{N}$) and abnormal outputs ($S_{NJ}$) in well-aligned models. Well-aligned models have safety neurons to reject malicious queries [51], but attack queries often bypass these and activate neurons linked to abnormal behaviors. As a result, normal and abnormal behaviors activate different neurons, causing coverage differences. It’s intriguing to observe that this phenomenon also occurs in non-aligned models. This may be because LLMs associate specific neurons with different knowledge domains [52, 20], leading to varying activation patterns depending on the query type. Abnormal behaviors, such as bias or harmful content, activate distinct knowledge areas, triggering neurons linked to negative or harmful knowledge domains, which results in noticeable activation differences. Note that in non-aligned models, the higher $S_{NJ}$ compared to $S_{NM}$ can be due to attack queries often involving unusual templates, instructions, or more diverse outputs, unlike straightforward malicious queries.
@@
Answer to RQ1: Considering the three requirements, NC and TKNC are relatively effective. They accurately identify synonymous and attack queries while ensuring stability and generalization across models. NLC demonstrates competitive performance but exhibits variability across different models.
@@ IV-C RQ2: Analyzing Model Layer-wise Contributions to Coverage in LLMs
Attention Layer versus MLP Layer: We first evaluate the performance of Attention and MLP layers from an overall model perspective. We use NC and TKNC, identified as effective in RQ1, as our research criteria and calculate RCG based on the data in Table IV. For NC, the RCG values for Attention layers in the four models are 6.00%, 15.67%, 19.28%, and 7.35%, respectively. Correspondingly, the RCG values for MLP layers are 4.29%, 2.16%, 0.76%, and 0.89%. It is evident that under the same criteria and models, NC is more effective in Attention layers than in MLP layers.
Similarly, for TKNC, the RCG values for Attention layers in the four models are 6.10%, 11.65%, 10.46%, and 11.32%, while the MLP layers show values of 3.43%, 21.09%, 6.99%, and 3.67%. Except for Llama-2-7B-Chat, the effectiveness of Attention layers remains superior to that of MLP layers. Additionally, as shown in Table IV, the $S_{NJ}$ coverage growth rates are generally highest for all criteria in Attention layers, with one exception. In contrast, in MLP layers, higher $S_{NM}$ growth rates are observed in NC for the Llama-2-7B-Chat model, and in TFC and NLC for the Pythia-12B model.
@@
Figure 4: The RCG Results Based on NC and TKNC for Different Blocks of the Four Target LLMs: OPT-125M Contains 12 Blocks, Llama-2-7B-Chat Contains 32 Blocks, Pythia-12B Contains 36 Blocks, and Gemma-2-27B-it Contains 46 Blocks
@@ IV-D RQ3: Investigating Token-level Impacts on Coverage Analysis in LLMs
Figure 5: The RCG Results Calculated Based on NC and TKNC for Different Tokens in the Target LLMs: Starting from the Last Token of Each Query, Each Model Compares 10 Consecutive Tokens
To investigate RQ3, we expand on the previous experiments by generating additional tokens for each test suite and calculate the coverage over the next 10 consecutive tokens, starting from the last token of the query  [^1] For experimental stability, we select the next token with the highest probability.. We explore how different tokens impact coverage analysis, assuming they may lead to varying coverage behaviors and reveal more comprehensive insights into model behavior. Our key question is: “How does generating additional tokens affect diversity assessment among test suites in LLM testing?”
Our further goal is to identify the optimal token positions for LLM testing to reduce computational costs and achieve efficient testing results. These points are crucial for determining the most effective moments to measure coverage, thereby optimizing the balance between testing and resource expenditure. After generating each new token, we calculate the corresponding coverage rates for the model. This process allows us to examine how the coverage evolves as the model generates more tokens beyond the initial query. Similarly, we analyze the impact of these tokens on coverage analysis through the RCG quantified by NC and TKNC. The experimental results are shown in Figure 5.
@@
V Application
We explore the following three applications. First, we propose Real-time Jailbreak Detection, using activation features from coverage criteria to classify queries as normal or jailbreak, enabling systematic identification of high-risk interactions. Second, we explore Test Case Prioritization, leveraging coverage levels to identify high-priority cases and remove redundancies, improving testing efficiency. Third, we investigate Jailbreak Case Generation, where coverage-guided methods refine prompts to generate adversarial examples.
Our study highlights neuron coverage (NC) as an effective criterion, with attention layers and the last query token showing higher sensitivity to test suites. These findings form the basis of our application-focused studies. More details are available on our website [46].
V-A Jailbreak Detection
TABLE V: Comparison of Jailbreak Detection Accuracy Across Models and Methods
Model Method Dataset Avg
Alpaca-gpt4 JailBreakV TruthfulQA GCG DeepInception Masterkey
OPT-125M Our 97.00% 91.40% 95.47% 88.50% 100.00% 99.00% 95.23%
Perplexity-Sentence 98.63% 8.97% 99.76% 67.00% 0.00% 0.00% 45.73%
Perplexity-Window 100.00% 0.07% 100.00% 5.00% 0.00% 0.00% 34.18%
PARDEN-Sentence 83.57% 46.13% 90.58% 31.00% 0.00% 21.00% 45.38%
PARDEN-Window 93.40% 37.63% 96.33% 27.50% 0.00% 2.50% 42.89%
Self-Reminder N/A 19.50% N/A 2.50% 5.00% 18.00% 11.25%
Cluster 100.00% 74.40% 100.00% 2.00% 0.00% 70.00% 57.73%
Llama-2-7B-Chat Our 95.60% 97.00% 91.43% 86.00% 100.00% 93.50% 93.92%
Perplexity-Sentence 98.70% 9.43% 97.55% 81.50% 0.00% 0.00% 47.86%
Perplexity-Window 97.83% 10.03% 99.39% 78.00% 0.00% 0.00% 47.54%
PARDEN-Sentence 57.37% 54.60% 38.92% 86.50% 100.00% 66.00% 67.23%
PARDEN-Window 54.70% 61.90% 34.52% 89.50% 98.00% 67.00% 67.60%
Self-Reminder N/A 60.50% N/A 82.00% 72.00% 80.00% 73.63%
Cluster 62.80% 98.00% 0.00% 100.00% 100.00% 100.00% 76.80%
Pythia-12B Our 94.60% 99.40% 82.13% 77.00% 100.00% 97.00% 91.69%
Perplexity-Sentence 97.20% 9.40% 98.78% 80.00% 0.00% 0.00% 47.56%
Perplexity-Window 100.00% 9.57% 100.00% 59.50% 0.00% 0.00% 44.85%
PARDEN-Sentence 85.20% 6.00% 83.60% 17.50% 34.00% 4.00% 38.38%
PARDEN-Window 85.00% 3.53% 81.64% 17.00% 36.00% 3.50% 37.78%
Self-Reminder N/A 18.50% N/A 5.50% 7.50% 11.00% 10.63%
Cluster 98.40% 54.40% 99.14% 0.00% 6.00% 12.00% 44.99%
Due to the significant differences in the parts of the model covered by normal queries and jailbreak attacks, we design a detection method based on the number of activated neurons (the feature used by NC). Specifically, we train an MLP model that uses the number of activated neurons as input to derive a binary result indicating whether a query triggers a jailbreak response. The input dimension is determined by the LLM’s architecture. For example, OPT-125M has 12 blocks, represented by a feature vector $(l_{0},l_{1},\ldots,l_{11})$, where each element is the activated neurons in the attention layer of a block. The classifier has four hidden layers with 256, 2048, 512, and 128 neurons. We validate its effectiveness on OPT-125M, Llama-2-7B-Chat, and Pythia-12B.
Dataset: We randomly select 2,500 queries from Alpaca-gpt4 as the training set and 500 as the test set, representing normal queries to LLMs. For JailBreakV-28k, we filter out all queries related to GCG (including attack suffixes) and select 2,500 attack queries as the training set and 500 as the test set, representing attack queries. From TruthfulQA, we select 817 queries as a validation set to evaluate the classifier’s generalization ability on normal queries. Additionally, we use 200 attack queries generated by GCG, 50 attack queries based on DeepInception [53], and 200 queries generated by Masterkey [10] as separate validation sets to evaluate the classifier’s generalization ability on attack queries.
@@ V-A Jailbreak Detection
Baseline: We select the most widely used perplexity filter [38], the state-of-the-art method PARDEN [40], and self-reminder [54] as baselines for comparison. Additionally, we design a jailbreak-attack detector based on the clustering method used in our empirical study for further comparison. Following [40], the perplexity filter’s threshold is set to the maximum perplexity of malicious queries in AdvBench, with a window size of 10. PARDEN uses a 0.6 threshold and a window size of 30. The clustering method uses the same training set to determine cluster centers, and jailbreak detection is performed by comparing the distances to these centers.
Evaluation results: As shown in Table V, our method achieves high average accuracies: 95.23% on OPT-125M, 93.92% on Llama-2-7B-Chat, and 91.69% on Pythia-12B, demonstrating its effectiveness across models.
The Perplexity-based filters exhibit excellent performance on normal queries (nearly 100% accuracy) but are unstable on attack queries. They perform well on GCG attacks, which increase perplexity through adversarial suffixes (e.g., 81.50% on Llama-2-7B-Chat), but struggle with jailbreak attacks like JailBreakV, DeepInception, and Masterkey that rely on templates and instructions. The performance of PARDEN varies across datasets and models. It performs well on normal queries for OPT-125M and Pythia-12B but poorly on attack queries (generally below 30% accuracy). Conversely, on Llama-2-7B-Chat, it achieves better results on attack queries (over 85% accuracy on multiple datasets) but underperforms on normal queries. The effectiveness of the Self-Reminder method depends on model alignment, achieving 73.63% on Llama-2-7B-Chat but only 10% on less aligned models.The Clustering method tends to be extreme in its classifications, performing well on the test set but poorly generalizing on some validation sets. It misclassifies validation queries on OPT-125M and Pythia-12B as normal, and those on Llama-2-7B-Chat as attacks, leading to inconsistent performance.
V-B Test Case Prioritization
TABLE VI: Accuracy of Threshold-Based Test Case Prioritization Classification